# Forward modeling fluctuations in the DESI LRGs target sample using image simulations

### v1

@@ 2.3 COSMOS Deep Catalogs
The COSMOS Deep Catalogs are a set of image products derived from DECam observations in the COSMOS region, processed with the same Legacypipe procedures as the main DR9 data. The only difference is that the main survey catalog has a maximum depth limit. The depth is closely related to the total number of individual exposures chosen for measurement in one processing unit. Thus, standard Legacy Surveys DR9 release uses far fewer individual images in the COSMOS regions than have been observed by DECam. The Deep catalogs, on the other hand, take as many available exposures as possible. Because of the large number of exposures used in the COSMOS Deep production run, the computational time used to process these data in the same footprint is much higher than that in the DR9 production run. For example, brick 1503p015 uses 113 hours of wall time in the COSMOS Deep production run on one CPU node on NERSC’s Cori machine, while in the DR9 production run it uses 8 hours of wall time on one node. The maximum numbers of overlapping exposures used in the DR9 production run in this region was 6, 7, 7 for $g,r,z$ bands, respectively. Meanwhile, the same numbers for the COSMOS deep production run are 45, 52, 120. This results in a much deeper depth in the three optical bands: The maximum psfdepth for the DR9 production run in this region are 25.19, 24.65, and 23.80 for $g$, $r$ and $z$, while the median depth are 22.99, 23.94, 24.47 for the same bands. The corresponding numbers for the COSMOS Deep production run are 26.53, 26.27, 25.73 for maximum depth, and 25.31, 25.53, 24.79 for median depth. Because of this depth, the COSMOS Deep Catalog reaches a higher precision in flux and shape measurement. It also detects more faint sources in the image.
The depth of the COSMOS Deep Catalog allows us to use these data as a “truth” catalog, and our injected targets are sampled from it. However, even within the COSMOS Deep Catalog, there are significant variations in the depth and we first determine a minimum depth cut for the truth sample. The Legacypipe works in such a way that the complexity of the light profile chosen to obtain the photometry of a given source depends highly on the depth, based on simple criteria on the $\chi^{2}$ improvement required to allow a more complex profile (shown in equation 5.1). We must make a trade-off in the truth catalog between its precision in shape, and its total galaxy counts. The “truth input” needs to be accurate enough in flux so that the output color distribution in simulated galaxies matches the real galaxies. Section 7.1 explains that the noise within the truth input leads to a mismatch in magnitude between real and synthetics LRGs. Figure 2 shows how these two variables compete given different depth cuts. We use the galdepth metric produced by Legacypipe, which gives the sensitivity to a round exponential-profile galaxy with a half-light radius 0.45". We determine the depth cut by checking the percentage of Sérsic galaxies in the sample. A higher fraction of Sérsic galaxies in the sample means that the galaxies are better resolved in morphology. The determination of the type of a galaxy in Legacypipe has a prior that favors a simpler type (further explained in section 5.1). Because of this prior, the images have to be clear enough for the pipeline to choose the type Sérsic for the galaxies in the images. By selecting regions with a higher fraction of Sérsic galaxies, we have a higher confidence in the truth sample’s LRG morphology, and the sample’s photometry will also be better. Based on the observed distribution of Sérsic fraction, we selected all data with galdepth_z $>24.5$ into our truth sample.
@@
The high depth in optical bands also affects the WISE band: the WISE band images remain the same as in the DR9 release, with a depth much lower than the optical band. The sources in the WISE images are measured in a process called “forced photometry”: We take the shape and location measured in the optical band, and fit these sources in the WISE images to obtain their flux values in the WISE bands. Having a higher depth in the optical band has both pros and cons for understanding the WISE measurements. It helps improve the accuracy in WISE band fitting by providing more precise information about the galaxies’ location and shape. However, since more faint sources are detected in optical band, the sources in the WISE images are more likely to be blended with other sources. In the presence of blending, the flux on one source could be falsely assigned to a nearby source due to a slight deviation in the input values of shape and PSF.
[…]
@@
As the COSMOS Deep data set is imaged with the DECam telescope, the filter properties are different from the ones used in the BASS/Mzls surveys. The COSMOS Deep Catalog is therefore not a representative truth sample for the BASS/MzLS region. Although it is potentially possible to map the truth catalog to the BASS/MzLS region with filter transmission function similar to what is described in equation 1-6 of [31], we were unable to perform a robust validation in limited tests. We therefore focus on the DECaLS data in this study and leave simulating BASS/MzLS for future work. However, given the similarities in the sample selection between the North and South imaging regions, our expectation is that any conclusions on the properties that drive fluctuations in target density based on our analysis of DECaLS data would hold for BASS/MzLS data.
@@
When sampling from the Deep catalogs, we debias the galaxy shapes (this will be discussed in section 5.1). We also apply the LRG-like cut (described in section 4), which reduces the sample to only the galaxies that have a reasonable likelihood of being selected as a DESI target. The final product yields 23,892 sources, with 4,986 that pass the DESI LRG target selection.
@@ 3 The Obiwan Pipeline
The synthetic source injection pipeline used in this work is Obiwan. It injects galaxies into the real images, and extracts them out using the same procedure that is applied to real galaxies in the Legacy Surveys DR9 data. This pipeline can be divided into two stages: In the first stage, it inject sources and processes images on optical bands, obtaining the information of flux, shape, etc for each source. In the second stage, it injects and processes images on the WISE bands. The location and shape for each source, which is obtained in the first stage, is used in this stage as input. The pipeline fits WISE images to obtain the WISE flux for each source. Figure 4 shows the workflow of this pipeline.
Figure 4: Workflow of Obiwan. The blue, green, red, and purple panel represents the $g$, $r$, $z$, $W1$/$W2$ band respectively. The black and white pattern represents the real galaxies, and the yellow pattern represents the simulated galaxies. The workflow shows how the simulated galaxies are added to real images on a per-exposure basis. Only patches with simulated galaxies are actually fitted. For the fitting of the WISE image, we also need extra information on galaxies from the DR9 data release.
@@ 3.1 Optical Bands
As mentioned before, Obiwan works by adding simulated sources to real images. The sources are added on a per-exposure basis, at the level of each CCD detector. The RA, Dec location of the injected source is converted into pixel detector units, the added source is convolved with the PSF at that location, and Poisson noise is applied afterwards. In the Legacypipe processing, these individual exposures are transformed into common brick coordinates. Then, the pipeline proceeds to the source detection stage, where sources are detected and segmented into “blobs”. A “blob” is defined as a set of contiguous pixels with signal-to-noise greater than 6 (when including all images in the optical bands), plus an 8-pixel margin. The blob is defined to be the same region in brick pixel space across all of the optical bands. Each blob will contain at least one source and the photometry is fit independently in each blob. A key aspect that makes Obiwan more efficient is that in the source fitting step, we only fit blobs that overlap the true input RA,Dec of our simulated sources. This helps reduce processing time, and each iteration takes 75% less time than fitting over the full image.
@@ 3.2 WISE bands
Compared to processing a whole image, Obiwan uses much less processing time in the optical band, thanks to only processing a small fraction of sources within the “blobs” that have injected galaxies in them. To preserve this computationally efficient structure while computing WISE flux values for Obiwan sources, we need some extra operations on the part of the pipeline that computes WISE flux for galaxies.
When producing the DR9 catalog, all blobs were processed in optical bands. The WISE band fitting takes all of the source information from optical bands, and uses the information to perform “forced photometry”. This means that the source location and shape parameters are held fixed and the WISE flux is the only free parameter. WISE forced photometry simultaneously fits all sources detected in optical bands over the entire brick of WISE images.
In our Obiwan simulation run, we do not have the optical-band information for all sources. Due to the computationally efficient structure applied in optical bands, we only have the sources within the same blob as simulated sources; sources outside the selected blobs are not processed in previous steps, thus we do not know the details of the sources’ shapes and locations in these non-processed blobs. However, these sources are processed and recorded in the DR9 catalogs. These non-processed blobs have identical conditions as the sources processed in DR9, so their properties are identical to what is recorded in DR9 as well. As a result, we can take these sources’ entries in DR9 as input for the WISE flux forced photometry measurement.
@@ 4 LRG-like target selection
We define a new LRG-like cut that is used to perform target selection on the sample from the COSMOS Deep Catalog. The selected sample with this LRG-like cut served as the truth sample that was injected into DESI Legacy Imaging Survey DR9 images.
@@
The sources near selection boundaries can be scattered inside or outside of the LRG target selection boundary. To fully simulate the scattering feature of LRGs, we need to define a LRG-like color cut, which spans a larger color selection box than the LRG SV3 selection. At the same time, it also needs to be small enough to not spend too much computing time on obtaining the photometry of sources that are unlikely to ever have measured photometry that matches the LRG SV3 selection. We describe our procedure to generate such a color selection box as follows.
@@
We need to develop a selection scheme that contains LRGs both inside and outside the LRG SV3 color/magnitude selection due to photometric scattering. We denote this selection scheme LRG-like color cut. We use Potential-LRGs to quantify the fraction of LRGs inside the LRG-like cut: For a given LRG-like cut, we define scatter rate as the amount of Potential-LRGs outside LRG-like cut, divided by the total number of Potential-LRGs. We also define contamination rate as the number of sources inside this color cut, divided by the number of Potential-LRGs inside the LRG-like cut. Our goal for the LRG-like cut is to make the selected sample have a low scatter rate and a low contamination rate.
The following process describes our method to develop the LRG-like cut: We start from the LRG SV3 selection cut, which is composed of 4 equations described in equation 4.14.4 excluding bold numbers. We iteratively extend each selection boundary from equations 4.14.4 by 0.01 magnitudes, and compute the scatter rate for each equation. We choose the selection cut that corresponds to the lowest scatter rate, and update this selection cut to the new boundary (i.e., increased by 0.01 magnitudes). This would be a new data point in figure 5. We perform this process iteratively, widening the selection cuts little by little. Each step is a new realization of a color cut, and corresponds to a different scatter rate and contamination rate. Figure 5 shows this variation: Each point represents one potential LRG-like color cut. As we extend the selection boundaries, we have a lower scatter rate and a greater contamination rate.
We select the color cut with the lowest scatter rate, as ideally, our simulation would include all sources potentially selected as LRGs. Our final selection corresponds to a scatter rate of 0.009, and a contamination rate of 4.03. These numbers imply that we will include more than 99% of potential LRG sources and that 20% of the injected sources will ultimately be selected as LRGs. Follow-up of the less than 1% of the Potential LRGs not included in this cut reveals that they are either blended or mismatched and their photometry is too far away from LRG SV3 color cut. The existence of these Potential LRGs is not due to photometric scattering so it is not necessary to include them in our simulation.
Figure 5: The change in scatter rate and contamination rate as we extend the color selection boundaries. Each point represents a set of color cuts that define the potential LRG-like sample. From the right side to the left side of the plot, we go to a more extended color cut. It would include more potential LRGs, as well as other non-LRGs. We eventually chose the left-most point as our final color cut. This color cut is wide enough to fully mimic the scattering of galaxies outside and inside the LRG SV3 selection boundary while keeping the ratio of input LRGs in Obiwan simulations high enough.
@@
Here the bold number shows the amount of boundary moved in each color cut. The non-bold part is the original LRG-SV3 selection cut. Eventually, 23892 sources are selected as truth input.
@@ 5.1 Truth Data Generation
The first term is summed over pixels of a galaxy stamp. We have pixelized data for the observed image (image), the modeled image (model), and a map that indicates the variance for each pixel (inverse_variance). The second term is a prior that is a function of galaxy ellipticity. The third term increases with the complexity of the galaxy model, and is used to determine the favorable model for a given galaxy. The prior term imposed by ellipticity and the number of parameters is independent of the depth of the image. This means that at higher depth, this prior term becomes less significant since the inverse_variance term increases and the first term gets bigger than the rest. The ellipticity_prior term also increases with the intrinsic ellipticity of a galaxy. For galaxies with low inverse_variance, this helps avoid fitting background noise into the galaxy. However, the measured shape will be biased. This bias also exists in the COSMOS Deep catalog. If we take the shape information as it is, and construct galaxy stamps directly from the COSMOS Deep catalog, the bias in shape would be applied twice: First, the input catalog is biased in shape. Second, our pipeline fitting is the same as Legacypipe, so we would use the same ellipticity prior during the fitting. Our output result made with this method would have a different distribution in shape compared with the real LRGs. We would not be able to fully represent the LRG distribution morphologically.
We develop a procedure to de-bias the shape distribution in the COSMOS Deep truth distribution: We compute the coadd observed image and the coadd inverse_variance image, as well as the average PSF (point spread function) image of the COSMOS Deep data in each brick. Then we replace the coadd observed image with Gaussian noise based on pixel values from the inverse_variance image. Next, we add one simulated COSMOS Deep source to its original pixel location, with the shape and flux information taken from the COSMOS Deep dataset. We define a blob that has this source centered in the middle of the image, with a size of $128\times 128$ pixels. With these modifications, it is much faster for the pipeline to locate the designated blob, and fit the one galaxy that we injected. We then fit this blob using the same process in the Legacypipe, and Legacypipe returns the measured shape information. After we obtain the shape measured from Legacypipe, we compare it with the input shape. When they don’t match, we modify the input shape and run the pipeline for more time. We repeat this process until the output shape reaches the desired precision. Although there is only a prior on ellipticity in equation 5.1, we find that the parameters of galaxy radius and Sérsic index are also biased when we check the shape distribution of Obiwan galaxies that directly samples from the COSMOS Deep catalog. It is likely because parameters for galaxy shapes are correlated with each other. We define a quantity A that includes 3 shape parameters: $e$ (ellipticity), $shape\_r$ (half-light radius), and $sersic\_n$ (Sérsic index). It is defined by
$\begin{split}\textbf{A}_{i}&=(\texttt{sersic\_n}_{i}^{O}-\texttt{sersic\_n}_{\texttt{s}})^{2}/5.5+(\texttt{e}_{i}^{O}-\texttt{e}_{\texttt{s}})^{2}\\ &\quad+(\texttt{shape\_r}_{i}^{O}-\texttt{shape\_r}_{\texttt{s}})^{2}\end{split}$ (5.2)
[…]
The $X$ here denotes one of the parameters from $sersic\_n$, $e$ or $shape\_r$. The upper index "I" denotes the input value from ith iteration. The step is a constant number used to determine how much the new parameter is modified. The three parameters are all updated during each iteration. We added some constraints on the ellipticity $e$. We force galaxies with $e=0$ (round galaxies or point source) in the deep photometry remains $e=0$. Galaxies with non-zero ellipticity preserve their orientation.
These new set of parameters are accepted if
$\textbf{A}_{i+1}<\textbf{A}_{i}$ (5.4)
Otherwise, we make the $step$ size smaller and try again, with a maximum of 3 times per trial. The $step$ value is defined as 0.2, 0.1, 0.05 for 0 fail, 1 fail, and 2 fails. The $step$ value is chosen for the code to reach the correct value quickly and precisely.
There are many advantages to a procedure like this. First, it makes debiasing computationally possible by working on the coadd images instead of individual exposures. A COSMOS Deep brick takes more than 40 hours to finish on one NERSC’s Cori nodes. Since we need to iteratively fit on the same spot for each source involved, it would take even more time than a COSMOS Deep run. Second, it automatically takes into account the bias from blending, without actually being contaminated by nearby sources: The blended region is more noisy and has a lower value on inverse_variance image. Also, when blending exists in a pixel, it would be brighter and has a higher Poisson noise.
We test our shape-modified set on the COSMOS Repeats set. We inject these shape-modified galaxies into COSMOS Repeats images and recover the output shape with Obiwan. We compare the shape distribution of our simulated LRGs with the real LRGs in the same set and perform the Kolmogorov–Smirnov (KS) test over the distribution; the results are shown in Fig 6. One can observe that based on the $p$-values of the KS test, the distribution of the shape-modified catalog is much more similar to the DESI LRGs when compared with the un-modified version. The $p$-values of the un-modified data are near zero, thus clearly rejecting the possibility that the un-modified data are representative of DESI LRGs. Conversely, the $p$-values of the modified data increase in all cases and are acceptable in most, and thus the debiasing process yields considerable improvement in the output shape distribution of Obiwan LRGs.
Figure 6: Histogram comparison between the simulated LRGs and the real LRGs in the same COSMOS Repeats footprint. The dot and star curves in this plot represent different types of simulated galaxies. The dot dashed curve uses input seeds from the de-biased sample, and the star solid curve uses input seeds from the un-modified sample. The x-axis denotes the 10 sets of the COSMOS Repeats numbered 0 to 9, and the y-axis uses a p-value to indicate how close the histograms are between the simulated and the real galaxy shape distribution, using the KS test; 1 being the most similar and 0 being the least similar. These plots compare the shape parameters of ellipticity, Sérsic index, and the half-light radius.

### v2

@@ 2.3 COSMOS Deep Catalogs
The COSMOS Deep Catalogs are image products derived from DECam observations in the COSMOS region, processed with the same Legacypipe procedures as the main DR9 data. The only difference is that the main survey catalog has a maximum depth limit. The depth is closely related to the total number of individual exposures in one processing unit chosen for measurement. Thus, standard Legacy Surveys DR9 release uses far fewer individual images in the COSMOS regions than have been observed by DECam. The Cosmos Deep catalog uses all individual images available and reaches a much higher depth in all optical bands. The higher depth enables the COSMOS deep catalog to reach a better precision in flux and shape measurement, and more detection of faint sources. Details for computational resources used are in Appendix A.2. In addition, we perform a cut of galdepth_z > 24.5 to ensure that the catalog is deep enough in $z$-band. Details about this cut are described in Appendix B.1.
@@
The high depth in optical bands also affects the WISE band: the WISE band images remain the same as in the DR9 release, with a depth much lower than the optical band. The sources in the WISE images are measured in a process called “forced photometry”: We take the shape and location measured in the optical band, and fit these sources in the WISE images to obtain their flux values in the WISE bands. A higher depth in the optical band has pros and cons for understanding the WISE measurements. It helps improve the accuracy of WISE band fitting by providing more precise information about the galaxies’ location and shape. However, since more faint sources are detected in the optical band, the sources in the WISE images are more likely to be blended with other sources. In the presence of blending, the flux on one source could be falsely assigned to a nearby source due to a slight deviation in the shape, PSF, and astrometric position input values. We further test flux accuracy in the WISE band and describe the results in Appendix B.2. We find no significant discrepancy in the COSMOS deep catalog for the WISE flux, so we use it as our ‘truth catalog’.
@@
As the COSMOS Deep data set is imaged with the DECam telescope, the filter properties differ from those used in the BASS/Mzls surveys. Therefore, the COSMOS Deep Catalog is not a representative truth sample for the BASS/MzLS region. Although it is potentially possible to map the truth catalog to the BASS/MzLS region with a filter-transmission function similar to what is described in equation 1-6 of [33], we were unable to perform a robust validation in limited tests. Thus, we focus on the DECaLS data in this study and leave simulating BASS/MzLS for future work. However, given the similarities in the sample selection between the North and South imaging regions, we expect that any conclusions on the properties that drive fluctuations in target density based on our analysis of DECaLS data would hold for BASS/MzLS data.
@@
When sampling from the Deep catalogs, we de-bias the galaxy shapes (this will be discussed in Section 5.1). We also apply the LRG-like cut (described in Section 4), which reduces the sample to only galaxies with a reasonable likelihood of being selected as a DESI target. The final product yields 23,892 sources, with 4,986 that pass the DESI LRG target selection.
@@ 3 The Obiwan Pipeline
The synthetic source injection pipeline used in this work is Obiwan. It injects galaxies into the real images, and extracts them using the same procedure applied to real galaxies in the Legacy Surveys DR9 data. This pipeline can be divided into two stages: In the first stage, it injects sources and processes images on optical bands, obtaining the information on flux, shape, etc, for each source. In the second stage, it injects and processes images on the WISE bands. The location and shape of each source, obtained in the first stage, is used in this stage as input. The pipeline fits WISE images to get the WISE flux for each source. Figure 2 shows the workflow of this pipeline.
Figure 2: Workflow of Obiwan. The blue, green, red, and purple panel represents the $g$, $r$, $z$, $W1$/$W2$ band respectively. The black and white pattern represents the real galaxies, and the yellow pattern represents the simulated galaxies. The workflow shows how the simulated galaxies are added to real images on a per-exposure basis. Only patches with simulated galaxies are fitted. For the fitting of the WISE image, we also need extra information on galaxies from the DR9 data release.
@@ 3.1 Optical Bands
Obiwan works by adding simulated sources to real images. The sources are added on a per-exposure basis, at the level of each CCD detector. The RA, Dec location of the injected source is converted into pixel detector units, the added source is convolved with the PSF at that location, and Poisson noise is applied afterward. In the Legacypipe processing, these individual exposures are transformed into common brick coordinates. Then, the pipeline proceeds to the source detection stage, where sources are detected and segmented into “blobs”. A “blob” is defined as a set of contiguous pixels with signal-to-noise greater than 6 (when including all images in the optical bands), plus an 8-pixel margin. The blob is defined to be the same region in brick pixel space across all of the optical bands. Each blob will contain at least one source and the photometry is fit independently in each blob. A key aspect that makes Obiwan more efficient is that in the source fitting step, we only fit blobs that overlap the true input RA,Dec of our simulated sources. This helps reduce processing time, and each iteration takes 75% less time than fitting over the full image.
@@ 3.2 WISE bands
Compared to processing a whole image, Obiwan uses much less processing time in the optical band, thanks to only processing a small fraction of sources within the “blobs” that have injected galaxies in them. To preserve this structure while computing WISE flux values for Obiwan sources, we need some extra operations on the part of the pipeline that computes WISE flux for galaxies.
When producing the DR9 catalog, all blobs were processed in optical bands. The WISE band fitting takes all of the source information from optical bands, and uses the information to perform “forced photometry”. This means that the source location and shape parameters are held fixed and the WISE flux is the only free parameter. “Forced photometry” simultaneously fits all sources detected in optical bands over the entire brick of WISE images.
In our Obiwan simulation run, we do not have the optical-band information for all sources. Due to the computationally efficient structure applied in optical bands, we only have the sources within the same blob as simulated sources; sources outside the selected blobs are not processed in previous steps; thus we do not know the details of the sources’ shapes and locations in these non-processed blobs. However, these sources are processed and recorded in the DR9 catalogs. These non-processed blobs have conditions identical to those of the sources processed in DR9, so their properties are identical to those recorded in DR9. As a result, we can take these sources’ entries in DR9 as input for the WISE flux forced photometry measurement.
@@ 4 LRG-like target selection
We define a new LRG-like cut used to perform target selection on the sample from the COSMOS Deep Catalog. The selected sample with this LRG-like cut served as the truth sample injected into DESI Legacy Imaging Survey DR9 images.
@@
The sources near selection boundaries can be scattered inside or outside of the LRG target selection boundary. To fully simulate the scattering feature of LRGs, we need to define a LRG-like color cut, which spans a larger color selection box than the LRG SV3 selection. At the same time, it also needs to be small enough not to spend too much computing time obtaining the photometry of sources that are unlikely to have measured photometry that matches the LRG SV3 selection. We describe our procedure to generate such a color selection box as follows.
@@
Due to photometric scattering, we need to develop a selection scheme that contains LRGs both inside and outside the LRG SV3 color/magnitude selection. We denote this selection scheme LRG-like color cut. We use Potential-LRGs to quantify the fraction of LRGs inside the LRG-like cut: For a given LRG-like cut, we define scatter rate as the amount of Potential-LRGs outside LRG-like cut, divided by the total number of Potential-LRGs. We also define contamination rate as the number of sources inside this color cut, divided by the number of Potential-LRGs inside the LRG-like cut. Our goal for the LRG-like cut is to make the selected sample have a low scatter and contamination rates.
The following process describes our method to develop the LRG-like cut: We start from the LRG SV3 selection cut, which is composed of 4 equations described in equation 4.14.4 excluding bold numbers. We iteratively extend each selection boundary from equations 4.14.4 by 0.01 magnitudes, and compute the scatter rate for each equation. We choose the selection cut that corresponds to the lowest scatter rate, and update this selection cut to the new boundary (i.e., increased by 0.01 magnitudes). This would be a new data point in Figure 3. We perform this process iteratively, widening the selection cuts little by little. Each step is a new realization of a color cut, corresponding to different scatter and contamination rates. Figure 3 shows this variation: Each point represents one potential LRG-like color cut. As we extend the selection boundaries, we have a lower scatter rate and a greater contamination rate.
We select the color cut with the lowest scatter rate, as ideally, our simulation would include all sources potentially selected as LRGs. Our final selection corresponds to a scatter rate of 0.009, and a contamination rate of 4.03. These numbers imply that we will include more than 99% of potential LRG sources and that 20% of the injected sources will ultimately be selected as LRGs. Follow-up of less than 1%
Figure 3: The change in scatter and contamination rates as we extend the color selection boundaries. Each point represents a set of color cuts that define the potential LRG-like sample. From the right side to the left side of the plot, we go to a more extended color cut. It would include more potential LRGs, as well as other non-LRGs. We eventually chose the left-most point as our final color cut. This color cut is wide enough to fully mimic the scattering of galaxies outside and inside the LRG SV3 selection boundary while keeping the ratio of input LRGs in Obiwan simulations high enough.
@@
The bold number shows the number of boundaries moved in each color cut. The non-bold part is the original LRG-SV3 selection cut. Eventually, 23892 sources are selected as truth input.
@@ 5.1 Truth Data Generation
The first term is summed over pixels of a galaxy stamp. We have pixelized data for the observed image (image), the modeled image (model), and a map that indicates the variance for each pixel (inverse_variance). The second term is a prior that is a function of galaxy ellipticity. The third term increases with the complexity of the galaxy model, and is used to determine the favorable model for a given galaxy. The prior term imposed by ellipticity and the number of parameters is independent of the depth of the image. This means that at higher depth, this prior term becomes less significant since the inverse_variance term increases and the first term gets bigger than the rest. The ellipticity_prior term also increases with the intrinsic ellipticity of a galaxy. For galaxies with low inverse_variance, this helps avoid fitting background noise into the galaxy. However, the measured shape will be biased. This bias also exists in the COSMOS Deep catalog. If we take the shape information as it is, and construct galaxy stamps directly from the COSMOS Deep catalog, the bias in shape would be applied twice: First, the input catalog is biased in shape. Second, our pipeline fitting is the same as Legacypipe, so we would use the same ellipticity prior during the fitting. Our output result made with this method would have a different distribution in shape compared with the real LRGs. We would not be able to fully represent the LRG distribution morphologically. Thus, we develop a procedure to de-bias the shape distribution in the COSMOS Deep truth distribution. The procedure is described in Appendix C. This procedure modifies the input shape of truth catalogs. After this de-biasing procedure, the shape distribution of output Obiwan LRGs is consistent with DESI LRGs.